LOG-IN
Displaying reports 1-20 of 2612.Go to page 1 2 3 4 5 6 7 8 9 10 End
KAGRA MIR (Polarization)
Print this report.
MarcEisenmann - 21:08, Thursday 07 July 2022 (3021)Get code to link to this report
Comment to Measurement of losses as a function of mirror rotation (Click here to view original report: 3020)

Marc, Michael, Yuhang

Today we finished the cleaning of #5 and the 1.5 inch sapphire.

Then, we tried to fine tune the rotation angle of the mirror.

This time, we placed a flexible ruler on the mirror and could precisely (+/-0.2 mm) estimate the arc length of the rotation.

The results are attached in figure 1.

Actually I made a mistake and rotated the HWP by +0.7 deg before this measurement so we need to add this value to the effective rotation of the mirror.

Note that for some points the mirror was a bit tilted which creates stronger coupling between s and p polarizations.

That's why we repeated measurements at some rotation angle.

We started measurement at -2.9 deg rotation with HWP = 0 deg.

In the end, before starting polarization measurement, we checked the HWP position that minimized the Ip power to be +0.8 deg.

Images attached to this comment
3021_20220707140929_smallrots2ploss.png
KAGRA MIR (Polarization)
Print this report.
MarcEisenmann - 20:56, Thursday 07 July 2022 (3020)Get code to link to this report
Measurement of losses as a function of mirror rotation

Marc, Yuhang

As for any waveplate, theta should have a 90 degrees periodicity with maybe one of the 2 orthogonal axis of better quality than the other.

We performed large rotation of the mirror while injecting s polarization and results are reported in the attached figure.

We estimated the rotation angle by measuring the arc length between each rotation. The errorbar comes from uncertainty in this estimation (+/-5 mm).

We measured the Ip and Is intensities at the center of the mirror.

We found out that we have, as expected, a 90 degrees periodicity in the minimum of Ip (or equivalently the s to p polarization losses).

Furthermore, it seems that a +90 degrees rotation (+ means clockwise rotation from the laser side), gave better results than around 0 degrees.

We started measurement at this position that are attached in the following figures.

They show the only small improvement on the losses but this might be due to the remaining theta.

The mean theta at the center position is -2.09 deg.

Images attached to this report
3020_20220707135150_s2plossroll.png 3020_20220707135439_rollpol0.png 3020_20220707135442_rollpol30.png 3020_20220707135447_rollpol60.png 3020_20220707135455_rolllosspol30.png 3020_20220707135459_rolllosspol60.png 3020_20220707135544_rolldntheta.png
Comments related to this report
MarcEisenmann - 21:08, Thursday 07 July 2022 (3021)

Marc, Michael, Yuhang

Today we finished the cleaning of #5 and the 1.5 inch sapphire.

Then, we tried to fine tune the rotation angle of the mirror.

This time, we placed a flexible ruler on the mirror and could precisely (+/-0.2 mm) estimate the arc length of the rotation.

The results are attached in figure 1.

Actually I made a mistake and rotated the HWP by +0.7 deg before this measurement so we need to add this value to the effective rotation of the mirror.

Note that for some points the mirror was a bit tilted which creates stronger coupling between s and p polarizations.

That's why we repeated measurements at some rotation angle.

We started measurement at -2.9 deg rotation with HWP = 0 deg.

In the end, before starting polarization measurement, we checked the HWP position that minimized the Ip power to be +0.8 deg.

KAGRA MIR (Polarization)
Print this report.
MarcEisenmann - 18:59, Thursday 07 July 2022 (3019)Get code to link to this report
Comment to birefringence measurement of AZTEC #3 and estimation of losses (Click here to view original report: 3013)

There is a little discrepancy between the 2 ways we compute the s to p polarization losses.

One possible explanation could be that we were using the mean of theta and delta n from 7 measurements during about 5 days.

It is quite probable that alignment condition or translation stage position slightly drifted between each measurements making this mean value a bit different with some sharp features.

I attach to this entry the comparison between the direct estimation of losses and the delta n and theta computed for each input polarization angle.

There is no sharp feature anymore and we have better agreement between the 2 estimations !

Images attached to this comment
3019_20220707115844_losspol15.png 3019_20220707115847_losspol30.png 3019_20220707115852_losspol45.png 3019_20220707115857_losspol60.png 3019_20220707115900_losspol75.png
R&D (FilterCavity)
Print this report.
MichaelPage - 17:47, Thursday 07 July 2022 (3018)Get code to link to this report
Comment to OPO replacement - green generation and characterisation (Click here to view original report: 3008)

Here is a set of data points measured for generated green power versus thermistor resistance. The PDH lock was optimised at each temperature change. Note that only about 200 µW of infrared is sent to the OPO. The gap was just due to a note taking error, so we will fill out the rest of the points soon.

Images attached to this comment
3018_20220707104732_opogreentemp220707.jpg
R&D (FilterCavity)
Print this report.
YuhangZhao - 17:31, Thursday 07 July 2022 (3016)Get code to link to this report
Comment to OPO replacement - green generation and characterisation (Click here to view original report: 3008)

Yuhang and Michael

After the OPO alignment optimization, it was found that PDH signal was lost. Today, we found what is causing this problem: the OPO transmission PD is misaligned (drop from more than 5V to ~50mV). After aligning OPO transmission PD back, the PDH signal is recovered. (It was checked that the signal sent to EOM is 2Vrms, according to logbook2469, the modulation strength is 0.1-0.3rad/V, so we should have a modulation of 0.2-0.6rad. If we have 0.6rad, we should see about 10% power drop from carrier. If we have 0.2rad, the power goes to sidebands should be negligible. the modulation should be almost negligible since we didn't observe a power decrease after sending modulation)

Then we measured OPO generated green power as a function of OPO temperature. The measurement offset power on power meter is 3.3uW. We take measurement from temperature controller reading 7.7kOhm to 6.5kOhm with a step of 0.02kOhm.

These are the measurements results

7.700 7.680 7.659 7.640 7.620 7.601 7.579 7.560 7.540 7.519 7.500 7.480 7.459 7.440 7.420 7.400 7.380 7.361 7.339 7.320 7.299 7.280 7.260 7.240 7.220 7.200 7.179 7.160 7.140 7.119 7.099 7.080 7.059 7.040                                                    
93.0 93.4 94.0 94.3 94.5 94.6 94.6 93.8 94.2 93.7 94.6 96.0 97.7 99.8 102.2 105.3 109.0 112.5 116.8 120.5 122.8 123.0 120.0 114.5 106.9 99.0 92.5 87.0 81.7 77.6 73.0 68.7 65.0 62.6                                                    

However, during this measurement, we found that the OPO lock was having some issues due to the disturbance from temperature change while OPO kept locked.

Thus we performed again the measurements with a OPO unlock/lock between each measurement. This new measurement didn't have locking issue and will be reported later as a comment to this logbook.

R&D (FilterCavity)
Print this report.
YuhangZhao - 19:06, Wednesday 06 July 2022 (3015)Get code to link to this report
Trial to see how CC field affect squeezing measurement

Aritomi and Yuhang

We followed the procedure in elog1188 to do a squeezing zero span measurement. This time, spectrum analyzer KEYSIGHT N9320B is used to perform this measurement.

Center frequency = 100kHz, RBW=VBW=1kHz, sweep time = 2s, SR560 gain = 100, signal sent to green phase shifter is 1.25Hz, 0.7Vpp.

We made measurement with and without CC field going to OPO, but CC PLL was not locked. The green power was 25mW sent to OPO. The nonlinear gain was measured to be around 3. The result is shown in attached figure.

From this measurement, it seems CC is not contributing any difference in this case. The measurement is too noisy anyway.

Images attached to this report
3015_20220706120458_figure1.png
R&D (FilterCavity)
Print this report.
NaokiAritomi - 18:35, Wednesday 06 July 2022 (3014)Get code to link to this report
Nonlinear gain on 20220706
green power (mW) 0 25
OPO temperature (kOhm) 7.13 7.13
p pol PLL frequency (MHz) 230 180
BAB maximum (mV) 464 1460
Nonlinear gain 1 3.1
KAGRA MIR (Polarization)
Print this report.
MarcEisenmann - 12:12, Wednesday 06 July 2022 (3013)Get code to link to this report
birefringence measurement of AZTEC #3 and estimation of losses

estimation of birefringence

By combining the several polarization measurements of AZTEC #3, it is possible to compute its birefringence parameters (delta n and theta) as shown in figure 1.

I modified also a bit this analysis as follow :

Because we are only sensitive to the modulus of the birefringence parameters, when theta is negative I take its opposite to only have positive theta.

Also, because delta n is proportional to arcsin( I_po * sin(2 * theta) ^2 ) where I_po is the p polarization when injecting s polarization, there could be points on the mirror where the arcsin is not defined (eg its parameters larger than 1).

In that case, I express delta n as pi/2 + arcsin( I_po * sin(2 * theta) ^2  mod(1) ).

I also show in figures 2 and 3 the stress coefficients.

Interestingly, the folding/discontinuity in theta happens for large stress area.

estimation of losses

From the birefringence parameters, it is possible to compute the s to p polarization losses as sin(2*theta)^2 * sin(pi*d*delta n / lambda) with d = 0.155 m the mirror thickness and lambda the wavelength.

This losses should actually corresponds exactly to the p polarization power when injecting s polarization.

These 2 measurements are shown in the top row of figure 4 (the black circle show the beam area when installed in KAGRA). They match really well except in the area with theta folding/discontinuity. We are currently investigating how to combine these 2 measurements to smoothen theta.

Also, we computed the mean losses as follow :

  from direct Ip measurement from birefringence measurement
accross all mirror 0.79 % 0.95%
weighted by the beam power distribution 0.76 % 0.96%
inside ITM beam diameter 0.52 % 0.72 %
Images attached to this report
3013_20220706041919_dntheta.png 3013_20220706041924_s0.png 3013_20220706041928_s1.png 3013_20220706051204_s2ploss.png
Comments related to this report
MarcEisenmann - 18:59, Thursday 07 July 2022 (3019)

There is a little discrepancy between the 2 ways we compute the s to p polarization losses.

One possible explanation could be that we were using the mean of theta and delta n from 7 measurements during about 5 days.

It is quite probable that alignment condition or translation stage position slightly drifted between each measurements making this mean value a bit different with some sharp features.

I attach to this entry the comparison between the direct estimation of losses and the delta n and theta computed for each input polarization angle.

There is no sharp feature anymore and we have better agreement between the 2 estimations !

KAGRA MIR (Polarization)
Print this report.
MarcEisenmann - 19:02, Tuesday 05 July 2022 (3012)Get code to link to this report
polarization measurements of AZTEC #3

This entry reports polarization measurements of AZTEC #3 with polarization angle varying from 0 deg ( s polarization at the input) to 75 deg with 15 deg increments.

As expected from the new calibration (see 3007), the sum of the calibrated Is and Ip is constant and equal to 1.

Images attached to this report
3012_20220705120210_pol0.png 3012_20220705120215_pol15.png 3012_20220705120219_pol30.png 3012_20220705120223_pol45.png 3012_20220705120227_pol60.png 3012_20220705120231_pol75.png
KAGRA MIR (Scattering)
Print this report.
MarcEisenmann - 18:38, Tuesday 05 July 2022 (3011)Get code to link to this report
Comment to Setup of NIR camera (Click here to view original report: 3009)

Aritomi, Marc, Yuhang

We installed the camera on the PCI setup.

We changed its IP address, installed the Pylon software on the FC pc and can now access it.

Tuning a bit the gamma correction, exposure time and gain of the camera we could not see any IR beam on the mirror.

We will try again later on with a not moving mirror.

R&D (FilterCavity)
Print this report.
YuhangZhao - 20:41, Monday 04 July 2022 (3010)Get code to link to this report
Comment to OPO replacement - green generation and characterisation (Click here to view original report: 3008)

Yuhang and Michael

It was found that after flipping OPO, the alignment of OPO got worse and was hard to recover.

Today we checked again the alignment of OPO. We conclude that it is certainly that the flipping of OPO makes alignment worse. This is because that the mode matching condition is very much different depending on which side the laser enters OPO.

It was always questioned how we can make sure a good internal alignment. In fact, the alignment procedure of OPO is not so much different from our filter cavity. The only difference is that we cannot control the angle of OPO reflection surfaces as easily as the filter cavity mirrors.

For OPO input surface, we adjust injection beam to make injection and reflection overlap. But at the same time, we need to make sure the crystal transmission has a good shape.

For OPO end surface, we adjust input-coupler surface position to find interference.

At this step, the alignment is not optimized but we need to fix the input-coupler. Then we need to optimized mode matching. When optimizing mm, we need to pay attention to which lens is more sensitive. Pay attention also that after moving a lens, we basically just need to move the tilt of one steering mirror to recover alignment and compare the new mode matching HOM to understand if the mode matching is getting better or worse. 

After optimizing mode matching, if there is still misalignment. (It was found that misalignment level is different when optimizing mode matching) We need to adjust input-coupling mirror.

The attached two figures show two situations: 1. after optimizing mode matching, before final alignment of input coupling mirror 2. after final alignment of input coupling mirror

There is still space to further optimize OPO alignment. Note that the OPO transmission power is increased in Fig.2.

Images attached to this comment
3010_20220704134030_wechatimage20220704204010.jpg 3010_20220704134036_wechatimage20220704204018.jpg
KAGRA MIR (Scattering)
Print this report.
MarcEisenmann - 21:36, Friday 01 July 2022 (3009)Get code to link to this report
Setup of NIR camera

In order to measure scattering at 1064 nm, I borrowed the camera foreseen to replace the camera in transmission of the END mirror (bassler acA-2040-25gmNIR)

I found the camera but could not find its power supply nor its lens so I used the POE ethernet connection and a lens of a spare/broken (?) camera nearby the first target.

I connected it to the new dgs switch for the camera server + ethernet connection for the front end and data concentrator pc.

The camera got really hot so I checked the power delivery of the POE which seemed fine (2W provided below the 3.1 W limit).

To made this check I connected my laptop to the switch and tweaked a bit the switch password and IP address.

However, because my windows is in japanese I could not tune too much the requirement on 'jumbo packet' so we will have to finalize this with japanese people.

I set a static IP address for the camera and could see the camera image on pylon software.

However, it seems that the 1064nm efficiency is really low so I'm not sure how well we will be able to see scattering.

I will do this test when AZTEC #3 birefringence measurement is finished.

Comments related to this report
MarcEisenmann - 18:38, Tuesday 05 July 2022 (3011)

Aritomi, Marc, Yuhang

We installed the camera on the PCI setup.

We changed its IP address, installed the Pylon software on the FC pc and can now access it.

Tuning a bit the gamma correction, exposure time and gain of the camera we could not see any IR beam on the mirror.

We will try again later on with a not moving mirror.

R&D (FilterCavity)
Print this report.
MichaelPage - 20:34, Friday 01 July 2022 (3008)Get code to link to this report
OPO replacement - green generation and characterisation

Yuhang and Michael

We set up the OPO in the reversed configuration (incoupler mirror on the side of detection power meter) and generated green from the OPO. Then, the OPO was locked, and we attempted to measure the generated green power as a function of crystal temperature to find the optimal phase matching working point.

However, we didn't get very much power. Looking over a range of thermistor resistances 5 - 7.5 kOhm, we expected to see a large peak of generated green power somewhere, but the most we got was 42 uW on the detection power meter (Yuhang thesis fig 4.26 shows 110 mW of green near 7.3 kOhm resistance and a smaller 83 mW peak near 6.7 kOhm). Before the measurement we removed some ND filters before the input FI, and early in the measurement we removed an unnecessary beam splitter in the laser path then realigned the OPO. But, it looks like the internal alignment is quite bad now, compared to 3006, probably just from removing those optics.

Images attached to this report
3008_20220701133417_opogreen20220701.png
Comments related to this report
YuhangZhao - 20:41, Monday 04 July 2022 (3010)

Yuhang and Michael

It was found that after flipping OPO, the alignment of OPO got worse and was hard to recover.

Today we checked again the alignment of OPO. We conclude that it is certainly that the flipping of OPO makes alignment worse. This is because that the mode matching condition is very much different depending on which side the laser enters OPO.

It was always questioned how we can make sure a good internal alignment. In fact, the alignment procedure of OPO is not so much different from our filter cavity. The only difference is that we cannot control the angle of OPO reflection surfaces as easily as the filter cavity mirrors.

For OPO input surface, we adjust injection beam to make injection and reflection overlap. But at the same time, we need to make sure the crystal transmission has a good shape.

For OPO end surface, we adjust input-coupler surface position to find interference.

At this step, the alignment is not optimized but we need to fix the input-coupler. Then we need to optimized mode matching. When optimizing mm, we need to pay attention to which lens is more sensitive. Pay attention also that after moving a lens, we basically just need to move the tilt of one steering mirror to recover alignment and compare the new mode matching HOM to understand if the mode matching is getting better or worse. 

After optimizing mode matching, if there is still misalignment. (It was found that misalignment level is different when optimizing mode matching) We need to adjust input-coupling mirror.

The attached two figures show two situations: 1. after optimizing mode matching, before final alignment of input coupling mirror 2. after final alignment of input coupling mirror

There is still space to further optimize OPO alignment. Note that the OPO transmission power is increased in Fig.2.

YuhangZhao - 17:31, Thursday 07 July 2022 (3016)

Yuhang and Michael

After the OPO alignment optimization, it was found that PDH signal was lost. Today, we found what is causing this problem: the OPO transmission PD is misaligned (drop from more than 5V to ~50mV). After aligning OPO transmission PD back, the PDH signal is recovered. (It was checked that the signal sent to EOM is 2Vrms, according to logbook2469, the modulation strength is 0.1-0.3rad/V, so we should have a modulation of 0.2-0.6rad. If we have 0.6rad, we should see about 10% power drop from carrier. If we have 0.2rad, the power goes to sidebands should be negligible. the modulation should be almost negligible since we didn't observe a power decrease after sending modulation)

Then we measured OPO generated green power as a function of OPO temperature. The measurement offset power on power meter is 3.3uW. We take measurement from temperature controller reading 7.7kOhm to 6.5kOhm with a step of 0.02kOhm.

These are the measurements results

7.700 7.680 7.659 7.640 7.620 7.601 7.579 7.560 7.540 7.519 7.500 7.480 7.459 7.440 7.420 7.400 7.380 7.361 7.339 7.320 7.299 7.280 7.260 7.240 7.220 7.200 7.179 7.160 7.140 7.119 7.099 7.080 7.059 7.040                                                    
93.0 93.4 94.0 94.3 94.5 94.6 94.6 93.8 94.2 93.7 94.6 96.0 97.7 99.8 102.2 105.3 109.0 112.5 116.8 120.5 122.8 123.0 120.0 114.5 106.9 99.0 92.5 87.0 81.7 77.6 73.0 68.7 65.0 62.6                                                    

However, during this measurement, we found that the OPO lock was having some issues due to the disturbance from temperature change while OPO kept locked.

Thus we performed again the measurements with a OPO unlock/lock between each measurement. This new measurement didn't have locking issue and will be reported later as a comment to this logbook.

MichaelPage - 17:47, Thursday 07 July 2022 (3018)

Here is a set of data points measured for generated green power versus thermistor resistance. The PDH lock was optimised at each temperature change. Note that only about 200 µW of infrared is sent to the OPO. The gap was just due to a note taking error, so we will fill out the rest of the points soon.

KAGRA MIR (Scattering)
Print this report.
MarcEisenmann - 20:56, Thursday 30 June 2022 (3007)Get code to link to this report
calibration for AZTEC #3 measurement

Previous analysis had some biased due to the calibration of our signals.The new calibration should take into account non ideal PBS and differents photodetectors.

We can express our 2 signals as :

DC = (Rs * Is + Rp * Ip ) * p(t) * Ks

AC = ( (1 - Rs) * Is + (1-Rp) * Ip ) * p(t) * Kp

where

Rs and Rp are PBS reflectivities for s and p polarizations

Ks and Kp are gain of the photodetector for respective s and p polarization readout

p(t) time-varying power fluctuations.

By measuring all these parameters, we can then reconstruct Is and Ip powers from our lockin amplifiers signal.

After carefully aligning the QWP and HWP to have linear s polarization, we measured AC and DC signals for 10 mn while injecting first s then p polarizations.

The measurements are attached in figure 1.

From these measurements, I could compute all the coefficients as shown in figure 2.

Note that here p1 are power fluctuations while injecting s polarization and p2 the one when injecting p polarization.

Using the mean of each of these coefficients, I show in figure 3 Is and Ip, the respective s and p polarizations when injecting s polarization.

As expected, Is = 1 and Ip = 0 and are independent on power fluctuations as well.

Rs and Rp coefficients are compatible with their specifications.

 

Also, during this calibration we found that the old lockin amplifier readout seems noisier than the new one.

Especially, we could barely see the power fluctuations that were hidden inside the noise (at that time we had about 2 nW of power reaching the photodetector but still a factor 50 above dark noise)

 

Following this calibration we started the AZTEC #3 birefringence measurement.

Images attached to this report
3007_20220630135416_acdccalibration.png 3007_20220630135423_calibparameters.png 3007_20220630135754_isipspol.png
R&D (FilterCavity)
Print this report.
YuhangZhao - 11:45, Thursday 30 June 2022 (3006)Get code to link to this report
The new OPO in ATC is realigned and ready for more characterizations

Yuhang and Michael

We realigned the new OPO in ATC yesterday. The screws holding OPO input-coupling mirror are adjusted to maximize the alignment. After this adjustment, the first higher order modes are totally removed. This indicates that the OPO internal alignment is finally optimized.

The second order mode is still very high, but it looks to be a Lagurre mode. So this should be optimized by a better mode matching. It was checked that this mode is not a wrong polarization.

In ATC, we have a temperature controller. We can start to look for the best temperature for green production. In addition, when green beam can be found, we should also characterize its beam profile.

Images attached to this report
3006_20220630044506_wechatimage20220630114333.jpg 3006_20220630044511_wechatimage20220630114339.jpg
R&D (FilterCavity)
Print this report.
NaokiAritomi - 20:24, Wednesday 29 June 2022 (3005)Get code to link to this report
OPO threshold is much larger than before

[Aritomi, Yuhang, Michael]

To investigate the issue of low nonlinear gain, we measured nonlinear gain with different green power as shown in the attached figure. The current OPO threshold is 161 mW which is much larger than last year in elog2577. We optimized PD position for BAB transmission and GRMC alignment, but the nonlinear gain is still low.

Images attached to this report
3005_20220629132358_nonlineargain20220628.png
R&D (FilterCavity)
Print this report.
YuhangZhao - 00:52, Wednesday 29 June 2022 (3004)Get code to link to this report
Comment to Characterization of CMRR for homodyne detector (Click here to view original report: 3000)

The CMRR was characterized by sending 10kHz 500mVpk-pk noise to laser intensity. To get the CMRR value, we compared two situation. One is when the LO is sent to only one eye of homodyne, but ND1 was used to make sure no saturation. Therefore, if the ND1 is removed, the noise should be increased by 20dB. The other situation is as usual way to measure shot noise, which is to measure sub-DC spectrum. The result is shown in Fig.1.

Then we also took a shot noise measurement at low frequency. The shot noise becomes flat until 3Hz (Fig.2). This is reasonable considering that the LO RIN measured in logbook 2988 shows almost 40dB RIN increase from 20Hz to 1Hz. Although we didn't have a clear number of CMRR in the past, we know that we achieved flat shot noise until 10Hz. Compared with the old usual shot noise, we should need more than additional 40dB CMRR to achieve shot noise like Fig.2.

Personal note: the 80dB CMRR is actually limited by the noise we can send to main laser intensity. Remind that the intensity noise modulation channel has an efficiency of 0.1A/V. So when we send 0.5V, we are actually modulating by only 0.05A the laser current. I think we can at least increase this noise by a factor of 10. But I checked the manual of our laser and I haven't found what is the damage threshold for this channel. This needs to be confirmed.

Images attached to this comment
3004_20220628173828_cmrr.png 3004_20220628175126_shotnoise.png
KAGRA MIR (General)
Print this report.
MarcEisenmann - 20:39, Tuesday 28 June 2022 (3003)Get code to link to this report
Preparation of birefringence and TWE measurements

Abe, Marc, Yuhang

cleaning

We removed #4 from translation stage.

We inspected the optics with strong flashlight and green lights and the 'clean room' is now really dirty...

We decided to clean #3 before birefringence measurement (first surface cleaned and second should be dry tomorrow morning).

After the scattering measurement of #1, we applied first contact on one surface and will apply it on the second surface tomorrow before moving it to ATC for the TWE measurements.

measurement preparation

We installed the 2 razor blades on the translation stage and measured their position wrt to the last steering mirror on the injection bench (153 mm for vertical and 177 mm for horizontal).

We wanted to measure the beam position with laser diode current of 6 A and rotating the HWP to reduce the power but the beam shape became terrible (kind of 2 blobs)

We decreased the laser diode current to 1A and rotated the HWP to recover the same power as before.

The beam shape became reasonable again which means that we can proceed with briefringence measurements but need some investigation before being able to start absorption measurements.

We measured the beam position on the last steering mirror and tweaked the pitch/yaw of this mirror to have 0.049 deg AOI  in vertical and -0.072 deg AOI in horizontal (see attached figures).

Tomorrow we will do calibration and start birefringence measurements.

Images attached to this report
3003_20220628132226_20220628vertical.jpg 3003_20220628133817_20220628horizontal.jpg
R&D (FilterCavity)
Print this report.
NaokiAritomi - 23:24, Monday 27 June 2022 (3002)Get code to link to this report
Comment to BS pitch/yaw picomotors seem to be broken (Click here to view original report: 2962)

I found an old elog420 which also says that BS Y picomotor doesn't work.

R&D (FilterCavity)
Print this report.
NaokiAritomi - 20:22, Monday 27 June 2022 (3001)Get code to link to this report
Nonlinear gain is too low

[Aritomi, Yuhang]

We checked the nonlinear gain with 25mW green (MZ offset: 4.2), but it is only ~2. It was 6.8 one year ago in elog2577. We optimized green alignment to OPO, OPO temperature, and p pol PLL frequency, but the nonlinear gain is still around 2. Here is the summary of today's measurement.

green power (mW) 0 25 25 25 25
OPO temperature (kOhm) 7.166 7.166 7.156 7.146 7.136
p pol PLL (MHz) 295 255 235 220 200
BAB maximum (mV) 98.4 216 226 230 232
Nonlinear gain 1 2.2 2.3 2.3 2.4